Ruth Lynn Seltzer

80 Orchard Street, Apt. 2 Medford, MA 02155

(617) 395-8244

Education

Wellesley College Wellesley, MA Sept. 1986– June 1990

Bachelor of Arts Degree in Psychology. Minor in Computer Science.

Work Experience

Fall  Badger Engineers, Inc.  Cambridge, MA

1990–Present Worked in Badger’s Word Processing Department for three months. Project secretary for Chevron PAR Project start-up under Gary Luck. Current project secretary responsible for all secretarial aspects of the McNeil Specialty Products Company Project.

1987–1990  Massachusetts Institute of Technology  Cambridge, MA

Wrote manuals for staff training and introductory guides for faculty and student use. Supported computers used for Electrical Engineering and Computer Science Department courses and faculty.

Summer 1989 Creative Temporary Agency Plainview, NY Assignments completed:

Fairchild Weston – Compiled documention for computerized weapon

training systems.

NEC America – General office support. Lincoln Savings Bank – Mortgage Processing, word processing. Loral – Accounting, spreadsheet processing.

Jerome I. Silverman Company Plainview, NY {Summer 1989, \atop 1987}

Developed and documented database software to track incoming mail, sales and inventory.

Summer 1986 Syosset Central School District Syosset, NY Managed stage production of summer theatrical production.

Summer 1986 McDonalds Restaurant Syosset, NY Responded to customer needs and inquiries during peak hours. Handled cash transactions.

Skills

Familiar with Lotus 1-2-3, Allways, Dbase, DisplayWrite, Word Perfect and other word processing programs and databases. Able to read Spanish.

Activities

Designed posters for M.I.T. Lecture Series Committee. Competed at the 1989 National Unicycle Meet. Tutored for Computer Science at Wellesley College.